What is the equation of the line containing the point (6,2) and parallel to the line 3x+7y=70?
Answer by Emsley(35)

Rearranging your line get 7y=70-3x so y=10-3/7x So gradient= -3/7
The equation of a line through (x1,y1)with gradient m is
y-y1=m(x-x1) For this question x1=6 y1=2 and m=-3/7
So equation is y-2=-3/7(x-6)
Mulyiply both sides by 7 7y-14=-3x+18 So 7y+3x=32